# Quenched pair breaking by interlayer correlations as a key to superconductivity in La3Ni2O7

### v4

@@ Abstract
The recent discovery of superconductivity in bilayer nickelate La3Ni2O7 under high pressure [Fig. 1(a)] heralds a new class of high-$T_{\mathrm{c}}$ superconductors [1]. Without doping, this material exhibits superconductivity under pressure exceeding $14$ GPa with maximal critical temperature of $T_{\mathrm{c}}\simeq 80$ K [1, 2, 3, 4, 5]. Notable feature in La3Ni2O7 is a multiorbital nature of low-lying states already at the level of density functional theory (DFT) [1, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 10, 16, 17, 18, 19, 20] [Fig. 1(b)]. Namely, three electrons per unit cell are distributed over Ni-$e_{g}$ orbitals in the top and bottom square-planar lattices, whereas Ni-$t_{2g}$ orbitals are fully occupied, thereby inactive for the low-energy physics. The two layers are coupled dominantly via interlayer nearest-neighbor hopping (or hybridization) between Ni-$d_{z^{2}}$ orbitals ($t_{\perp}^{z}\simeq-0.63$ eV) [6, 8, 13]. The hopping between Ni-$d_{x^{2}-y^{2}}$ is much smaller ($|t_{\perp}^{x}|<0.05$ eV) [6, 8]. Most importantly, $t_{\perp}^{z}$ is deemed to be crucial for the noninteracting Fermi surface (FS) topology and theories of superconductivity in La3Ni2O7 [18, 21, 16, 13, 22, 23, 24, 25, 26, 12, 27, 28, 29, 19].
@@
In this paper, we employ a cluster (cellular) dynamical mean-field theory (CDMFT) [30, 31, 32] to address nonperturbatively the nonlocal as well as the local electronic correlations within the two-site clusters (dimers) of the bilayer square lattice model for La3Ni2O7 [Fig. 1(a)]. One of the key findings of our study is a Lifshitz transition resulting in a change of the FS topology which does not occur when only local correlations are taken into account. By solving an appropriate gap equation, we show that the IECs promote $s\pm$-wave pairing. The underlying mechanism is the quenching of ferromagnetic (FM) fluctuations resulting from the Lifshitz transition due to IECs. Based on this picture, we provide a possible reason of why superconductivity emerges only under high pressure.
@@
We consider a Hamiltonian on the bilayer square lattice: $\mathcal{H}=H_{0}+H_{\mathrm{int}}$. Here, $H_{0}$ is a tight-binding term for the Ni-$e_{g}$ subspace describing the band structure for which we use the maximally localized Wannier function (MLWF) description for the DFT result of La3Ni2O7 under high pressure (29.5 GPa) [8] [Fig. 1(b)]. $H_{\mathrm{int}}$ is the local interaction term between Ni-$e_{g}$ orbital electrons on the same Ni site, and is given by the standard Kanamori form consisting of $U$ (intraorbital Coulomb interaction), $J$ (Hund’s coupling), and $U^{\prime}$ (interorbital Coulomb interaction; $U^{\prime}=U-2J$). We use $U=3.7$ eV, $J=0.6$ eV, and $U^{\prime}=2.5$ eV by taking ab initio estimates for the $e_{g}$ MLWF model [34]. See Supplemental Material (SM) for more information [35].
@@
Here, ket symbols indicate the corresponding Wannier states with site index $i$ for the bilayer square lattice and spin $\sigma\in\{\uparrow,\downarrow\}$. $\bar{\eta}$ and $\underline{\eta}$ represent Ni-$e_{g}$ orbitals ($\eta\in\{x^{2}-y^{2},z^{2}\}$) in the top and bottom layers, respectively. Hereafter $x^{2}-y^{2}$ is denoted by $x$ and $z^{2}$ by $z$. Site and spin indices are omitted unless needed. In this BA basis the CDMFT self-energy becomes orbital-diagonal and momentum-independent, so our CDMFT is equivalent to “four-orbital single-site DMFT”. The interlayer hopping $t^{\eta}_{\perp}$ between $e_{g}$ orbitals $|d_{\bar{\eta}}\rangle$ and $|d_{\underline{\eta}}\rangle$ turns into a hybridization gap of $2|t^{\eta}_{\perp}|$ between BA orbitals $|\widetilde{d}_{\eta_{+}}\rangle$ and $|\widetilde{d}_{\eta_{-}}\rangle$. Thus a small (large) splitting is realized for $\eta=x$ ($\eta=z$) [schematically shown in the right panel of Fig. 1(b)].
@@
The IECs, however, significantly modify this picture [Fig. 2(b)]. While the $\alpha$ pocket remains nearly unchanged, the $\beta$ and the $\gamma$ pockets are largely affected by IECs. The $\beta$ pocket becomes more diamond-shaped with spectral weight at the first Brillouin zone (FBZ) boundary being shifted toward the $X$ point. We also find redistribution of electron occupations in favor of half-filled $\bar{z}$ and $\underline{z}$ orbitals with $\langle n_{\bar{z}}\rangle=\langle n_{\underline{z}}\rangle\simeq 0.93$ compared to the DMFT value of $0.85$. Most interestingly, the $\gamma$ pocket disappears which results in a Lifshitz transition of the FS. Looking at the orbital character of the FS [Fig. 2(c)] reveals that $x_{-}$ and $z_{-}$ (for the $\beta$ pocket around $X$ point) and $z_{+}$ (for the $\gamma$ pocket around the $M$ point) BA orbitals underlie the FS modification.
@@
In contrast to the $x_{\pm}$ components, large $t^{z}_{\perp}$ gives rise to strong IECs in the $z_{\pm}$ components. We investigate the real part $\Re\Sigma_{l}(i\omega_{n})$ which modifies the on-site energy level of the orbital $l$; see Fig. 2(e). See SM for the imaginary part [35]. We note first that the Hartree-Fock self-energy, $\Re\Sigma_{l}(i\omega_{\infty})$, does not modify the FS topology because $\Re\Sigma_{x_{\pm}}(i\omega_{\infty})-\Re\Sigma_{z_{\pm}}(i\omega_{\infty})$ is only about $0.1$ eV and $\Re\Sigma_{x_{+}/z_{+}}(i\omega_{\infty})=\Re\Sigma_{x_{-}/z_{-}}(i\omega_{\infty})$.
In a low-frequency regime ($\omega_{n}\ll 10$ eV), however, $\Re\Sigma_{z_{+}}(i\omega_{n})$ is smaller and $\Re\Sigma_{z_{-}}(i\omega_{n})$ is larger than the value at infinite frequency. This, in turn, shifts effectively the onsite energy levels of $z_{\pm}$ upward ($z_{-}$) and downward ($z_{+}$) with respect to their DFT counterparts, thereby enhancing the hybridization gap. In fact, this low-energy behavior is the origin of the shifts of spectral weight and concomitant FS change seen in Fig. 2(b). Near the $X$ point $z_{-}$ has substantial weight in the $\beta$ pocket of the noninteracting FS. Thus the large upturn of $\Re\Sigma_{z_{-}}(i\omega_{n})$ as $\omega_{n}\rightarrow 0$ makes an upward shift in energy near the $X$ point, leading to the change of the $\beta$ pocket in CDMFT; see also SM [35]. The physics here bears a close resemblance to that of VO2 in which intersite correlations within dimers promote intradimer singlets with an enhanced hybridization gap [41, 42, 43].
@@
A phase transition to the superconducting state occurs when the corresponding pairing susceptibility diverges, which requires numerical evaluation of the pairing vertex $\mathbf{\Gamma}^{\mathrm{s/t}}$ for singlet (s) or triplet (t) Cooper pairs [52, 53] (bold symbols will be used to denote vectors and matrices). The spin and charge susceptibilities ($\bm{\chi}^{\mathrm{sp/ch}}$) and the related irreducible vertices ($\mathbf{\Gamma}^{\mathrm{sp/ch}}$) contribute to $\mathbf{\Gamma}^{\mathrm{s/t}}$. Calculating frequency- and momentum-dependent $\mathbf{\Gamma}^{\mathrm{sp/ch}}$ and $\bm{\chi}^{\mathrm{sp/ch}}$, however, is highly nontrivial for multiorbital systems. We thus follow an idea previously employed to study cuprates, ruthenates, and iron-based superconductors [54, 55, 56, 57]. Namely, $\mathbf{\Gamma}^{\mathrm{sp/ch}}$ are parametrized by effective intraorbital Coulomb interaction $\overline{U}$ and Hund’s coupling $\overline{J}$, i.e., $\mathbf{\Gamma}^{\mathrm{sp/ch}}\rightarrow\overline{\mathbf{\Gamma}}^{\mathrm{sp/ch}}(\overline{U},\overline{J})$ (we assume the interorbital value $\overline{U}^{\prime}=\overline{U}-2\overline{J}$). The effective vertices $\overline{\mathbf{\Gamma}}^{\mathrm{sp/ch}}(\overline{U},\overline{J})$ are independent of frequency and momentum, see SM [35]. This leads to the gap equation
@@
We find the predominance of singlet over triplet pairings arising from antiferromagnetic (AFM) fluctuations. Figure 3(a) presents the resulting superconducting phase diagram for the leading singlet channel. Since we cannot pinpoint the precise magnitude of $\overline{U}$ and $\overline{J}$, we scan a range of values. The vertical axis is given by the Stoner enhancement factor $\alpha_{\mathrm{sp}}$ which indicates the maximum eigenvalue of $\overline{\mathbf{\Gamma}}^{\mathrm{sp}}(\overline{U},\overline{J})\bm{\chi}^{0}(\bm{q},i\nu_{0})$ and gauges the proximity to a magnetic instability. Here, $\bm{\chi}^{0}$ is the irreducible susceptibility, $\chi^{0}_{lml^{\prime}m^{\prime}}(q)=-\frac{T}{N}\sum_{k}G_{ll^{\prime}}(k+q)G_{m^{\prime}m}(k)$, which is the lowest-order term of the spin susceptibility, $\bm{\chi}^{\mathrm{sp}}=\bm{\chi}^{0}[\mathbf{1}-\overline{\mathbf{\Gamma}}^{\mathrm{sp}}(\overline{U},\overline{J})\bm{\chi}^{0}]^{-1}$. Thus, $\alpha_{\mathrm{sp}}$ is determined entirely from $\overline{U}$ and $\overline{J}$, provided $\bm{\chi}^{0}(\bm{q},i\nu_{0})$ is given [35]. For both DMFT and CDMFT cases, the leading pairing symmetry in the $e_{g}$-orbital basis is always the intraorbital $s$-wave/interorbital $d_{x^{2}-y^{2}}$-wave pairing; see the inset of Fig. 3(a). Projecting to the noninteracting FS, this pairing corresponds to the $s\pm$-wave where the gap changes sign between the neighboring FS pockets [35]. This leading pairing symmetry is in line with many previous studies [21, 13, 12, 8, 23, 25, 18, 22]. The qualitative features of the gap functions remain unchanged over the entire parameter range while $\bm{Q}$ (the crystal momentum at which the maximum eigenvalue of $\overline{\mathbf{\Gamma}}^{\mathrm{sp}}(\overline{U},\overline{J})\bm{\chi}^{0}(\bm{q},i\nu_{0})$ emerges) changes from $\bm{Q}=(\pi,0)$ for $\overline{J}/\overline{U}\lessapprox 0.06$ to an incommensurate wave vector around the $M$ point for $\overline{J}/\overline{U}\gtrapprox 0.06$ for both DMFT and CDMFT as highlighted in Fig. 3(b); see SM for the details [35].
@@
To further corroborate this argument, we analyze how the DMFT superconducting instabilities are affected by the FM fluctuation by introducing a scaling factor $\zeta$ for $\chi^{0}_{z_{+}z_{+}z_{+}z_{+}}(q)$. Namely, $\chi^{0}_{z_{+}z_{+}z_{+}z_{+}}(q)$ is rescaled to $\zeta\chi^{0}_{z_{+}z_{+}z_{+}z_{+}}(q)$ before constructing $\bm{\chi}^{\mathrm{sp/ch}}$ and $\bm{\Gamma}^{\mathrm{s}}$. Indeed, as shown in the lower panel of Fig. 3(c), the interlayer FM spin susceptibility $\chi^{\mathrm{sp}}_{\bar{z}\bar{z}\underline{z}\underline{z}}(q=0)$ turns AFM with decreasing $\zeta$ followed by an attractive pairing interaction $\Gamma^{\mathrm{s}}_{\bar{z}\bar{z}\underline{z}\underline{z}}(q=0)$ and an increase of $\lambda_{\mathrm{sc}}$. See SM for additional data [35]. Note also that since the $\gamma$ pocket is the only FS pocket dispersive along the $k_{z}$ direction [19], the disappearance of the $\gamma$ pocket by IECs makes La3Ni2O7 effectively two-dimensional.
@@
We now turn to the question of “Why does superconductivity emerge only in the high-pressure phase?”. A useful insight is obtained from a recent experiment which reports that pressure mainly shrinks the out-of-plane Ni-O bond length while the in-plane one is weakly affected [58]. Thus, the main effect of pressure can be addressed with the change of $t^{z}_{\perp}$ which is sensitive to the out-of-plane Ni-O bond length. Since $t^{z}_{\perp}\simeq-0.63$ eV at 29.5 GPa under which superconductivity emerges [6, 8, 13], a smaller magnitude of $t^{z}_{\perp}$ should correspond to the lower pressure case. In light of this observation, we investigate two “low pressure” cases, namely $t^{z}_{\perp}=-0.45$ eV and $t^{z}_{\perp}=-0.55$ eV, using CDMFT.
@@
In Fig. 4(a), we find that large $\chi^{0}_{z_{+}z_{+}z_{+}z_{+}}(\bm{q},i\nu_{0})$ emerges for the two low-pressure cases, in sharp contrast to the high-pressure result ($t^{z}_{\perp}\simeq-0.63$ eV) which we have already noticed in Fig. 3. Interestingly, this result provide a plausible scenario of why La3Ni2O7 is not superconducting in the low-pressure phase because FM $\chi^{0}_{z_{+}z_{+}z_{+}z_{+}}(\bm{q},i\nu_{0})$ obstructs the singlet pairing as detailed above. While the $\gamma$ pocket gives rise to strong FM fluctuations, the actual magnetic transition occurs at a finite $\bm{q}$ as shown in Fig. 4(b) which presents the maximum eigenvalue of the spin susceptibility at each $\bm{q}$, $\chi^{\mathrm{sp}}_{\mathrm{max}}(\bm{q},i\nu_{0})$, for $t^{z}_{\perp}=-0.45$ eV. $\chi^{\mathrm{sp}}_{\mathrm{max}}(\bm{q},i\nu_{0})$ shows a peak at $\bm{q}=\bm{Q}_{\mathrm{SDW}}$ which is different from but close to the SDW wave vector $\bm{Q}^{\mathrm{exp.}}_{\mathrm{SDW}}=(0.5,0.5)\pi$ reported by experiments [51, 50]. See SM for further discussion [35].
@@
Figure 4: (a) $\chi^{0}_{z_{+}z_{+}z_{+}z_{+}}(\bm{q},i\nu_{0})$ calculated using CDMFT Green’s functions for three different values of $t^{z}_{\perp}$ (in units of eV). (b) $\chi^{\mathrm{sp}}_{\mathrm{max}}(\bm{q},i\nu_{0})$ at $\alpha_{\mathrm{sp}}=0.95$ and $\overline{J}/\overline{U}=0.2$ for $t^{z}_{\perp}=-0.45$ $\mathrm{eV}$. (c) FSs obtained from DMFT and CDMFT for the same values of $t^{z}_{\perp}$. We used $U=3.7$ eV, $J=0.6$ eV, and $U^{\prime}=U-2J$ for all the cases. $\bm{Q}_{\mathrm{SDW}}$ in (b) is highlighted with the orange arrow.
@@
To trace the origin of the difference between the two low-pressure cases and the high-pressure ($t^{z}_{\perp}\simeq-0.63$ eV) case, we investigate the CDMFT FSs [lower panels in Fig. 4(c)]. The strength of IECs is controlled by $|t^{z}_{\perp}|$, so the shape of the $\beta$ and the $\gamma$ pockets in CDMFT FS is basically the same as the DMFT FS for the smallest $|t^{z}_{\perp}|$ ($t^{z}_{\perp}=-0.45$ eV). As $|t^{z}_{\perp}|$ increases, however, the $\gamma$ pocket gets suppressed in CDMFT which is consistent with the evolution of $\chi^{0}_{z_{+}z_{+}z_{+}z_{+}}(\bm{q},i\nu_{0})$ presented in Fig. 4(a); see SM for further discussion on the microscopic origin, especially on the change of the $\beta$ pocket [35]. At $t^{z}_{\perp}\simeq-0.63$ eV, the FS which we have already seen in Fig. 2(b) is realized. In contrast, the FS is almost unaffected by $t^{z}_{\perp}$ in DMFT. Hence, it can be seen that the concerted effect of pressure (as modelled via $t^{z}_{\perp}$) and IECs induces the Lifshitz transition. This transition quenches the FM channel resulting in an enhancement of the singlet pairing mediated by AFM fluctuations.

### v5

@@ Abstract
The recent discovery of superconductivity in bilayer nickelate La3Ni2O7 under high pressure [Fig. 1(a)] heralds a new class of high-$T_{\mathrm{c}}$ superconductors [1]. Without doping, this material exhibits superconductivity under pressure exceeding $14$ GPa with maximal critical temperature of $T_{\mathrm{c}}\simeq 80$ K [1, 2, 3, 4, 5]. A notable feature in La3Ni2O7 is a multiorbital nature of low-lying states already at the level of density functional theory (DFT) [1, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 10, 16, 17, 18, 19, 20] [Fig. 1(b)]. Namely, three electrons per unit cell are distributed over Ni-$e_{g}$ orbitals in the top and bottom square-planar lattices, whereas Ni-$t_{2g}$ orbitals are fully occupied, thereby inactive for the low-energy physics. The two layers are coupled dominantly via interlayer nearest-neighbor hopping (or hybridization) between Ni-$d_{z^{2}}$ orbitals ($t_{\perp}^{z}\simeq-0.63$ eV) [6, 8, 13]. The hopping between Ni-$d_{x^{2}-y^{2}}$ is much smaller ($|t_{\perp}^{x}|<0.05$ eV) [6, 8]. Most importantly, $t_{\perp}^{z}$ is deemed to be crucial for the noninteracting Fermi surface (FS) topology and theories of superconductivity in La3Ni2O7 [18, 21, 16, 13, 22, 23, 24, 25, 26, 12, 27, 28, 29, 19].
@@
In this Letter, we employ a cluster (cellular) dynamical mean-field theory (CDMFT) [30, 31, 32] to address nonperturbatively the nonlocal as well as the local electronic correlations within the two-site clusters (dimers) of the bilayer square lattice model for La3Ni2O7 [Fig. 1(a)]. One of the key findings of our study is a Lifshitz transition resulting in a change of the FS topology which does not occur when only local correlations are taken into account. By solving an appropriate gap equation, we show that the IECs promote $s\pm$-wave pairing. The underlying mechanism is the quenching of ferromagnetic (FM) fluctuations resulting from the Lifshitz transition due to IECs. Based on this picture, we provide a possible reason of why superconductivity emerges only under high pressure.
@@
We consider a Hamiltonian on the bilayer square lattice: $\mathcal{H}=H_{0}+H_{\mathrm{int}}$. Here, $H_{0}$ is a tight-binding term for the Ni-$e_{g}$ subspace describing the band structure for which we use the maximally localized Wannier function (MLWF) description for the DFT result of La3Ni2O7 under high pressure (29.5 GPa) [8] [Fig. 1(b)]. $H_{\mathrm{int}}$ is the local interaction term between Ni-$e_{g}$ orbital electrons on the same Ni site, and is given by the standard Kanamori form consisting of $U$ (intraorbital Coulomb interaction), $J$ (Hund’s coupling), and $U^{\prime}$ (interorbital Coulomb interaction; $U^{\prime}=U-2J$). We use $U=3.7$, $J=0.6$, and $U^{\prime}=2.5$ eV by taking ab initio estimates for the $e_{g}$ MLWF model [34]. See Supplemental Material for more information [35].
@@
Here, ket symbols indicate the corresponding Wannier states with site index $i$ for the bilayer square lattice and spin $\sigma\in\{\uparrow,\downarrow\}$. $\bar{\eta}$ and $\underline{\eta}$ represent Ni-$e_{g}$ orbitals ($\eta\in\{x^{2}-y^{2},z^{2}\}$) in the top and bottom layers, respectively. Hereafter $x^{2}-y^{2}$ is denoted by $x$ and $z^{2}$ by $z$. Site and spin indices are omitted unless needed. In this BA basis the CDMFT self-energy becomes orbital diagonal and momentum independent, so our CDMFT is equivalent to “four-orbital single-site DMFT”. The interlayer hopping $t^{\eta}_{\perp}$ between $e_{g}$ orbitals $|d_{\bar{\eta}}\rangle$ and $|d_{\underline{\eta}}\rangle$ turns into a hybridization gap of $2|t^{\eta}_{\perp}|$ between BA orbitals $|\widetilde{d}_{\eta_{+}}\rangle$ and $|\widetilde{d}_{\eta_{-}}\rangle$. Thus a small (large) splitting is realized for $\eta=x$ ($\eta=z$) [schematically shown in the right panel of Fig. 1(b)].
@@
The IECs, however, significantly modify this picture [Fig. 2(b)]. While the $\alpha$ pocket remains nearly unchanged, the $\beta$ and the $\gamma$ pockets are largely affected by IECs. The $\beta$ pocket becomes more diamond shaped with spectral weight at the first Brillouin zone (FBZ) boundary being shifted toward the $X$ point. We also find redistribution of electron occupations in favor of half-filled $\bar{z}$ and $\underline{z}$ orbitals with $\langle n_{\bar{z}}\rangle=\langle n_{\underline{z}}\rangle\simeq 0.93$ compared to the DMFT value of $0.85$. Most interestingly, the $\gamma$ pocket disappears which results in a Lifshitz transition of the FS. Looking at the orbital character of the FS [Fig. 2(c)] reveals that $x_{-}$ and $z_{-}$ (for the $\beta$ pocket around $X$ point) and $z_{+}$ (for the $\gamma$ pocket around the $M$ point) BA orbitals underlie the FS modification.
@@
In contrast to the $x_{\pm}$ components, large $t^{z}_{\perp}$ gives rise to strong IECs in the $z_{\pm}$ components. We investigate the real part $\Re\Sigma_{l}(i\omega_{n})$ which modifies the on-site energy level of the orbital $l$; see Fig. 2(e). See Supplemental Material for the imaginary part [35]. We note first that the Hartree-Fock self-energy, $\Re\Sigma_{l}(i\omega_{\infty})$, does not modify the FS topology because $\Re\Sigma_{x_{\pm}}(i\omega_{\infty})-\Re\Sigma_{z_{\pm}}(i\omega_{\infty})$ is only about $0.1$ eV and $\Re\Sigma_{x_{+}/z_{+}}(i\omega_{\infty})=\Re\Sigma_{x_{-}/z_{-}}(i\omega_{\infty})$.
In a low-frequency regime ($\omega_{n}\ll 10$ eV), however, $\Re\Sigma_{z_{+}}(i\omega_{n})$ is smaller and $\Re\Sigma_{z_{-}}(i\omega_{n})$ is larger than the value at infinite frequency. This, in turn, shifts effectively the onsite energy levels of $z_{\pm}$ upward ($z_{-}$) and downward ($z_{+}$) with respect to their DFT counterparts, thereby enhancing the hybridization gap. In fact, this low-energy behavior is the origin of the shifts of spectral weight and concomitant FS change seen in Fig. 2(b). Near the $X$ point $z_{-}$ has substantial weight in the $\beta$ pocket of the noninteracting FS. Thus the large upturn of $\Re\Sigma_{z_{-}}(i\omega_{n})$ as $\omega_{n}\rightarrow 0$ makes an upward shift in energy near the $X$ point, leading to the change of the $\beta$ pocket in CDMFT; see also Supplemental Material [35]. The physics here bears a close resemblance to that of VO2 in which intersite correlations within dimers promote intradimer singlets with an enhanced hybridization gap [41, 42, 43].
@@
A phase transition to the superconducting state occurs when the corresponding pairing susceptibility diverges, which requires numerical evaluation of the pairing vertex $\mathbf{\Gamma}^{\mathrm{s/t}}$ for singlet (s) or triplet (t) Cooper pairs [52, 53] (bold symbols will be used to denote vectors and matrices). The spin and charge susceptibilities ($\bm{\chi}^{\mathrm{sp/ch}}$) and the related irreducible vertices ($\mathbf{\Gamma}^{\mathrm{sp/ch}}$) contribute to $\mathbf{\Gamma}^{\mathrm{s/t}}$. Calculating frequency- and momentum-dependent $\mathbf{\Gamma}^{\mathrm{sp/ch}}$ and $\bm{\chi}^{\mathrm{sp/ch}}$, however, is highly nontrivial for multiorbital systems. We thus follow an idea previously employed to study cuprates, ruthenates, and iron-based superconductors [54, 55, 56, 57]. Namely, $\mathbf{\Gamma}^{\mathrm{sp/ch}}$ are parametrized by effective intraorbital Coulomb interaction $\overline{U}$ and Hund’s coupling $\overline{J}$, i.e., $\mathbf{\Gamma}^{\mathrm{sp/ch}}\rightarrow\overline{\mathbf{\Gamma}}^{\mathrm{sp/ch}}(\overline{U},\overline{J})$ (we assume the interorbital value $\overline{U}^{\prime}=\overline{U}-2\overline{J}$). The effective vertices $\overline{\mathbf{\Gamma}}^{\mathrm{sp/ch}}(\overline{U},\overline{J})$ are independent of frequency and momentum, see Supplemental Material [35]. This leads to the gap equation
@@
We find the predominance of singlet over triplet pairings arising from antiferromagnetic (AFM) fluctuations. Figure 3(a) presents the resulting superconducting phase diagram for the leading singlet channel. Since we cannot pinpoint the precise magnitude of $\overline{U}$ and $\overline{J}$, we scan a range of values. The vertical axis is given by the Stoner enhancement factor $\alpha_{\mathrm{sp}}$ which indicates the maximum eigenvalue of $\overline{\mathbf{\Gamma}}^{\mathrm{sp}}(\overline{U},\overline{J})\bm{\chi}^{0}(\bm{q},i\nu_{0})$ and gauges the proximity to a magnetic instability. Here, $\bm{\chi}^{0}$ is the irreducible susceptibility, $\chi^{0}_{lml^{\prime}m^{\prime}}(q)=-\frac{T}{N}\sum_{k}G_{ll^{\prime}}(k+q)G_{m^{\prime}m}(k)$, which is the lowest-order term of the spin susceptibility, $\bm{\chi}^{\mathrm{sp}}=\bm{\chi}^{0}[\mathbf{1}-\overline{\mathbf{\Gamma}}^{\mathrm{sp}}(\overline{U},\overline{J})\bm{\chi}^{0}]^{-1}$. Thus, $\alpha_{\mathrm{sp}}$ is determined entirely from $\overline{U}$ and $\overline{J}$, provided $\bm{\chi}^{0}(\bm{q},i\nu_{0})$ is given [35]. For both DMFT and CDMFT cases, the leading pairing symmetry in the $e_{g}$-orbital basis is always the intraorbital $s$-wave and interorbital $d_{x^{2}-y^{2}}$-wave pairing; see the inset of Fig. 3(a). Projecting to the noninteracting FS, this pairing corresponds to the $s\pm$-wave where the gap changes sign between the neighboring FS pockets [35]. This leading pairing symmetry is in line with many previous studies [21, 13, 12, 8, 23, 25, 18, 22]. The qualitative features of the gap functions remain unchanged over the entire parameter range while $\bm{Q}$ (the crystal momentum at which the maximum eigenvalue of $\overline{\mathbf{\Gamma}}^{\mathrm{sp}}(\overline{U},\overline{J})\bm{\chi}^{0}(\bm{q},i\nu_{0})$ emerges) changes from $\bm{Q}=(\pi,0)$ for $\overline{J}/\overline{U}\lessapprox 0.06$ to an incommensurate wave vector around the $M$ point for $\overline{J}/\overline{U}\gtrapprox 0.06$ for both DMFT and CDMFT as highlighted in Fig. 3(b); see Supplemental Material for the details [35].
@@
To further corroborate this argument, we analyze how the DMFT superconducting instabilities are affected by the FM fluctuation by introducing a scaling factor $\zeta$ for $\chi^{0}_{z_{+}z_{+}z_{+}z_{+}}(q)$. Namely, $\chi^{0}_{z_{+}z_{+}z_{+}z_{+}}(q)$ is rescaled to $\zeta\chi^{0}_{z_{+}z_{+}z_{+}z_{+}}(q)$ before constructing $\bm{\chi}^{\mathrm{sp/ch}}$ and $\bm{\Gamma}^{\mathrm{s}}$. Indeed, as shown in the lower panel of Fig. 3(c), the interlayer FM spin susceptibility $\chi^{\mathrm{sp}}_{\bar{z}\bar{z}\underline{z}\underline{z}}(q=0)$ turns AFM with decreasing $\zeta$ followed by an attractive pairing interaction $\Gamma^{\mathrm{s}}_{\bar{z}\bar{z}\underline{z}\underline{z}}(q=0)$ and an increase of $\lambda_{\mathrm{sc}}$. See Supplemental Material for additional data [35]. Note also that since the $\gamma$ pocket is the only FS pocket dispersive along the $k_{z}$ direction [19], the disappearance of the $\gamma$ pocket by IECs makes La3Ni2O7 effectively two dimensional.
@@
We now turn to the question of “Why does superconductivity emerge only in the high-pressure phase?” A useful insight is obtained from a recent experiment which reports that pressure mainly shrinks the out-of-plane Ni-O bond length while the in-plane one is weakly affected [58]. Thus, the main effect of pressure can be addressed with the change of $t^{z}_{\perp}$ which is sensitive to the out-of-plane Ni-O bond length. Since $t^{z}_{\perp}\simeq-0.63$ eV at 29.5 GPa under which superconductivity emerges [6, 8, 13], a smaller magnitude of $t^{z}_{\perp}$ should correspond to the lower pressure case. In light of this observation, we investigate two “low pressure” cases, namely $t^{z}_{\perp}=-0.45$ and $t^{z}_{\perp}=-0.55$ eV, using CDMFT.
@@
In Fig. 4(a), we find that large $\chi^{0}_{z_{+}z_{+}z_{+}z_{+}}(\bm{q},i\nu_{0})$ emerges for the two low-pressure cases, in sharp contrast to the high-pressure result ($t^{z}_{\perp}\simeq-0.63$ eV) which we have already noticed in Fig. 3. Interestingly, this result provides a plausible scenario of why La3Ni2O7 is not superconducting in the low-pressure phase because FM $\chi^{0}_{z_{+}z_{+}z_{+}z_{+}}(\bm{q},i\nu_{0})$ obstructs the singlet pairing as detailed above. While the $\gamma$ pocket gives rise to strong FM fluctuations, the actual magnetic transition occurs at a finite $\bm{q}$ as shown in Fig. 4(b) which presents the maximum eigenvalue of the spin susceptibility at each $\bm{q}$, $\chi^{\mathrm{sp}}_{\mathrm{max}}(\bm{q},i\nu_{0})$, for $t^{z}_{\perp}=-0.45$ eV. $\chi^{\mathrm{sp}}_{\mathrm{max}}(\bm{q},i\nu_{0})$ shows a peak at $\bm{q}=\bm{Q}_{\mathrm{SDW}}$ which is different from but close to the SDW wave vector $\bm{Q}^{\mathrm{exp.}}_{\mathrm{SDW}}=(0.5,0.5)\pi$ reported by experiments [51, 50]. See Supplemental Material for further discussion [35].
@@
Figure 4: (a) $\chi^{0}_{z_{+}z_{+}z_{+}z_{+}}(\bm{q},i\nu_{0})$ calculated using CDMFT Green’s functions for three different values of $t^{z}_{\perp}$ (in units of eV). (b) $\chi^{\mathrm{sp}}_{\mathrm{max}}(\bm{q},i\nu_{0})$ at $\alpha_{\mathrm{sp}}=0.95$ and $\overline{J}/\overline{U}=0.2$ for $t^{z}_{\perp}=-0.45$ $\mathrm{eV}$. (c) FSs obtained from DMFT and CDMFT for the same values of $t^{z}_{\perp}$. We used $U=3.7$, $J=0.6$ eV, and $U^{\prime}=U-2J$ for all the cases. $\bm{Q}_{\mathrm{SDW}}$ in (b) is highlighted with the orange arrow.
@@
To trace the origin of the difference between the two low-pressure cases and the high-pressure ($t^{z}_{\perp}\simeq-0.63$ eV) case, we investigate the CDMFT FSs [lower panels in Fig. 4(c)]. The strength of IECs is controlled by $|t^{z}_{\perp}|$, so the shape of the $\beta$ and the $\gamma$ pockets in CDMFT FS is basically the same as the DMFT FS for the smallest $|t^{z}_{\perp}|$ ($t^{z}_{\perp}=-0.45$ eV). As $|t^{z}_{\perp}|$ increases, however, the $\gamma$ pocket gets suppressed in CDMFT which is consistent with the evolution of $\chi^{0}_{z_{+}z_{+}z_{+}z_{+}}(\bm{q},i\nu_{0})$ presented in Fig. 4(a); see Supplemental Material for further discussion on the microscopic origin, especially on the change of the $\beta$ pocket [35]. At $t^{z}_{\perp}\simeq-0.63$ eV, the FS that we have already seen in Fig. 2(b) is realized. In contrast, the FS is almost unaffected by $t^{z}_{\perp}$ in DMFT. Hence, it can be seen that the concerted effect of pressure (as modeled via $t^{z}_{\perp}$) and IECs induces the Lifshitz transition. This transition quenches the FM channel resulting in an enhancement of the singlet pairing mediated by AFM fluctuations.